# Moving toward best practice when using propensity score weighting in survey observational studies

### v1

@@ 4.2.2 Is there a universal winner among the augmented estimators when all models are correctly specified?
Our simulation results, when both the propensity score and outcome models are correctly specified, indicate that all three augmented estimators—moment estimator (MOM), clever covariate estimator (CVR), and weighted regression estimator (WET)—consistently outperform conventional propensity score weighting estimators (PSW). While CVR demonstrate high efficiency under good overlap (Table 5), its sensitivity to different sampling mechanisms and overlap scenarios reduces its practical reliability. MOM offers consistent robustness but provides less bias control and coverage stability compared to WET across all estimands. Consequently, WET with W.PS emerges as the most reliable and efficient estimator, maintaining near-nominal coverage, lower bias, and superior performance, particularly under poor overlap (Table 6) and treatment-dependent sampling (Web Table 9 and 10).
@@
Focusing first on the ”Cor|Cor” columns under good overlap (Table 5), W.PS-based MOM, CVR, and WET outperform PSW across all estimands (PATE, PATT, and PATO). The CVR estimator achieves significant efficiency gains, notably for PATT and PATO. Meanwhile, WET consistently performs slightly better than MOM, maintaining around 93% coverage and biases below 2% across most rows.
@@
Under poor overlap (”Cor|Cor” columns in Table 6) PSW alone produces sizable biases (e.g., 8.848% for PATE) and reduced coverage (76.1%). This highlights the well-known vulnerability of simple weighting methods under poor overlap.[36] In contrast, MOM, CVR, and WET demonstrate lower biases and significantly improved efficiency and coverage. For example, CVR with W.PS achieves nearly twenty-fold efficiency gains for PATE, reaching a relative efficiency of 23.8. Similarly, CVR boosts efficiency for PATT and PATO while maintaining better control over coverage. However, comparing across the augmented estimators, CVR and MOM show greater sensitivity under poor overlap. For instance, CVR achieves only 80% coverage for PATE, and MOM drops to 57.8% coverage for PATT. In comparison, WET consistently maintains coverage near or above 90%, excelling particularly in PATO, where the inherent robustness of PATO protects all methods—both PSW and augmented estimators—enabling them to achieve nominal-level coverage. Among these, WET with W.PS stands out, achieving the highest nominal coverage for PATO at 94.2% and outperforming other estimators.
@@
When sampling probabilities depend on treatment assignment (Web Tables 9 and 10), the advantage of WET becomes even more evident. Under good overlap scenarios (Web Table 9), WET with W.PS achieves minimal biases and relative efficiency exceeding 1.5 for PATE and PATT while maintaining nominal coverage around 94%, demonstrating its resilience under favorable overlap conditions. By contrast, CVR suffers from higher bias and reduced coverage, particularly for PATT, where extreme survey weights correlated with the group variable destabilize its performance. Similar patterns emerge under poor overlap (Web Table 10), where WET maintains near-nominal coverage for PATO. MOM shows better stability than CVR under poor overlap but still falls short of WET’s efficiency.
@@
Taken together, WET with W.PS excels under complex sampling mechanisms. The weighted regression framework effectively minimizes bias and maintains stable coverage, even under extreme overlap conditions. While MOM provides consistent robustness and CVR achieves high efficiency under ideal conditions as in Table 5, both are more vulnerable to instability when survey weights are heterogeneous. Thus, WET with W.PS stands out as the most robust choice , particularly when the sampling mechanism depends on treatment and generates highly variable survey weights.
@@
4.2.3 Across all misspecification scenarios, which methods are more robust and should be recommended under the complex survey setting?
When model specifications deviate from the truth, the performance of estimators varies by estimand and overlap. Examining the “Mis|Cor,” “Cor|Mis,” and “Mis|Mis” columns in Tables 5 and 6 reveals distinct advantages of MOM and WET over CVR, particularly under poor overlap and misspecified outcome models.
@@
When the propensity score model is misspecified but the outcome model is correct under good overlap (“Mis|Cor” columns in Table 5), all three augmented estimators achieve nominal coverage for PATE, demonstrating their doubly robust properties compared to PSW alone, which struggles to achieve even 80% coverage. MOM and WET control bias and coverage better than CVR. In poor overlap scenarios (Table 6), CVR achieves smaller bias but suffers reduced coverage, dropping to only 80%. In contrast, WET with W.PS maintains robustness, achieving 94.1% coverage for PATO, even with misspecified propensity score models.
@@
When the propensity score model is correct but the outcome model is misspecified (“Cor|Mis” columns), CVR demonstrates significant instability. Its bias increases dramatically for PATE under good overlap (Table 5), reaching nearly 10% compared to 0.2% bias for MOM and WET, and its coverage drops below 50% under poor overlap (Table 6). This highlights CVR’s strong reliance on the correct outcome model to manage the extreme values introduced by clever covariates in poor overlap scenarios. While MOM struggles with PATT coverage, WET consistently performs best, achieving the optimal efficiency-coverage balance for PATO, despite the poor overlap and outcome model misspecification.
@@
Under the most challenging scenario, where both the propensity score and outcome models are misspecified (“Mis|Mis” columns), all methods experience significant biases and deteriorated coverage. Despite these challenges, WET remains the most reliable option, consistently maintaining coverage around 80–85% under good overlap, while CVR and MOM frequently see coverage drop below 80%. The extent of bias varies depending on the type of misspecification. In our simulation settings, where interaction terms are omitted from both models, this misspecification strongly affects propensity score estimation and outcome regression, leading to highly biased estimates and poor coverage across all estimators.
@@
In conclusion, among the augmented estimators, while the clever covariate estimator (CVR) demonstrates potential for high efficiency with correctly specified models under good overlap, its heightened sensitivity to model misspecification and poor overlap limits its practical reliability. In contrast, both the moment estimator (MOM) and weighted regression estimator (WET) exhibit greater robustness, with WET emerging as the most resilient choice across varying survey settings, maintaining low bias and near-nominal coverage for a wide range of estimands, even under challenging conditions. Thus, WET with W.PS is strongly recommended for reliable performance in real-world applications.
5 Case studies
@@ 5.1 Application to the ECLS-K dataset
In this first data example, we present a real-world example with multi-stage sampling to demonstrate the application of the different methods for using propensity score weighting with complex survey data. The data used are from the The Early Childhood Longitudinal Study, Kindergarten class 1998-1999 (ECLS-K), [37] a nationally representative longitudinal study that tracks early school experiences of a large cohort from kindergarten through eighth grade in the United States. The study collects comprehensive information at the child, household, and school levels, with survey weights reflecting the multi-stage probability sampling design employed to ensure representativeness. [16] The participants were selected prior to the assignment of treatment. Specifically, children were sampled at the start of the study upon their entry into kindergarten, and subsequent treatment, such as the receipt of special education services, was determined in later stages, typically based on administrative records collected during follow-up data collection waves, such as in first grade or fifth grade. This example reflects a prospective study design, consistent with our main simulation setting, in which the sampling process does not depend on treatment assignment and survey weights are explicitly known. This design simplifies the integration of survey weights into subsequent estimations by reducing the steps required. Corresponding estimators tailored for prospective studies are further detailed in the appendix. The treatment status was defined as participation in special education services, as documented in school administrative records. The outcome of interest was the IRT-scaled math achievement test score from the ECLS-K dataset. A set of 39 baseline covariates covering demographic, household, and school characteristics were included to address potential confounding, as described in Keller and Tipton.[38]
@@
Web Figure 1 illustrates the population-level covariate balance, measured by Population-level survey-weighted Standardized Mean Differences (PSMD), which align with the weights $(\omega_{1,p}(\mathbf{X}_{i}),\omega_{0,p}(\mathbf{X}_{i}))$ used in the effect estimation and correspond to the specifications outlined in Table 1 for each target population and estimand. The PSMD for a baseline covariate $X$ is calculated as the absolute difference between the weighted means of $X$ for the treated and control groups, scaled by the pooled weighted standard deviation $s_{p}$. Specifically, PSMD is defined as |$\textbar\bar{X}_{1}-\bar{X}_{0}\textbar/s_{p}$, where the weighted means are $\bar{X}_{1}=\left(\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})X_{i}Z_{i}\right)/\left(\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\right)$ for the treated group and $\bar{X}_{0}=\left(\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})X_{i}(1-Z_{i})\right)/\left(\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})\right)$ for the control group. The pooled weighted standard deviation, $s_{p}$, reflects a composite measure of the variability of $X$ within the treated and control groups, weighted according to the propensity score weights $\omega_{1,p}$ and $\omega_{0,p}$, and accounts for the sample size in each group. The figure compares the covariate balance for each population-level estimand—PATE, PATT, and PATO—using survey-weighted regression (W.PS). The visual evidence clearly demonstrates that W.PS achieves superior covariate balance, as indicated by a larger proportion of covariates achieving PSMDs below the 0.10 threshold, a criterion for adequate balance.[39] This is further detailed in the grey-shaded cells of Web Table 17 , which shows how W.PS systematically improves covariate balance across most baseline characteristics. Specifically, W.PS achieves exact balance for PATO, further supporting its appropriateness for estimands where balancing the overlap population is critical.
To estimate the causal effect of elementary school special education services on fifth-grade math achievement, we applied the Propensity Score Weighting (PSW) estimator alongside three augmented estimators: the Moment Estimator (MOM), Clever Covariate Regression (CVR), and Weighted Regression Estimator (WET). Building on theoretical results and simulation studies that consistently identify survey-weighted propensity scores (W.PS) as the most robust and reliable approach across diverse conditions, this analysis emphasizes the application of W.PS. The estimated treatment effects for PATE, PATT, and PATO using W.PS are presented in Table 7, underscoring its practical advantages in achieving reliable causal inference.
The estimated treatment effect of special education services on fifth-grade math achievement using the WET under W.PS for PATO was -5.80, with a 95% confidence interval of (-7.34, -4.26). This significant negative effect indicates that, on average, students in special education scored approximately 5.8 points lower in math compared to the general population with similar baseline characteristics. In contrast, the PSW estimator produced larger standard errors, such as SE = 2.80 for PATT, and confidence intervals that included zero, highlighting its lower precision and reduced ability to confirm significant effects. The augmented estimators (MOM, CVR, and WET) consistently yielded confidence intervals that excluded zero across all estimands, demonstrating sufficient statistical power for detecting the underlying causal effect. Among them, WET under W.PS consistently yielded more precise and robust estimates, whereas MOM and CVR exhibited greater variability, particularly under potential model misspecification in real-world analyses. Notably, the PATO estimand remained stable across all estimators, emphasizing the reliability of the overlap population for deriving comparable treatment effects. These results align with theoretical guarantees and simulation findings, further supporting the superior performance of WET when combined with W.PS. The evidence underscores WET’s practical utility and robustness in addressing complex survey designs and achieving reliable causal estimates.
Table 7: Estimated treatment effects and standard errors (SE) for PATE, PATT, and PATO using different estimators under W.PS in the ECLS-K study.
PSW MOM CVR WET PATE (SE) PATT (SE) PATO (SE)
-3.31(2.64) -1.12(2.80) -5.81(1.61)
-4.47(1.54) -3.68(1.90) -5.96(1.61)
-3.91(1.47) -4.38(1.28) -4.65(1.15)
-4.86(1.43) -4.58(1.64) -5.80(1.53)
@@ 5.2 Application to the 2009 Medical Expenditure Panel Survey dataset
In our second empirical example, we demonstrate how the proposed methods perform in a retrospective survey design where the sampling mechanism depends directly on the exposure—here, race. We use data from the 2009 Medical Expenditure Panel Survey (MEPS),[AHRQ2012] which historically oversamples minority groups, including Hispanics and Blacks, to improve the precision of health disparity analyses. Starting in 2006, this strategy expanded to include Asian households in response to changing demographics and analytical priorities.[40] This oversampling design ensures adequate representation of minority groups but can also lead to disproportionate survey weights, particularly for subgroups with limited population overlap. Unlike Data Example 1 (ECLS-K), where participants were sampled before treatment status was realized, MEPS employs a retrospective design where individuals are sampled after their race is determined, leading to survey weights that often correlate with race and demographic strata. This underscores the importance of incorporating these survey weights into propensity score and effect estimation.
@@
The 2009 MEPS public-release file includes 9830 non-Hispanic White adults (hereafter “Whites”), 4020 Blacks, 1446 Asians, and 5150 Hispanics, each assigned a survey weight reflecting their probability of selection and nonresponse adjustments.[41, 40] Here, we focus on two pairwise comparisons (White–Asian and White–Hispanic) to estimate racial disparities in total annual healthcare expenditures, measured in 2009 USD. Although “race” is not a manipulable treatment, the methodological goal parallels standard propensity score approaches to balancing covariates between two groups. Following the framework for unconfounded descriptive comparisons described in Li and Li,[42] we control for a set of variables reflecting clinical need, as instructed by the Institute of Medicine (IOM),[43, 44] including body mass index (BMI), SF-12 physical and mental component summaries, health conditions, age, gender, and marital status. Additionally, we account for selected socioeconomic and functional limitation variables, ensuring alignment with established approaches to health care disparity analyses. These variable selections align with prior applications of MEPS data in disparity research,[41, 36, 45, 42] while recognizing that some SES variables, such as education and region, are not fully incorporated due to limitations in the available dataset.
@@
Web Figures 2, 3, and 4 collectively illustrate the interplay between survey weights, propensity score distributions estimated by the survey-weighted propensity score model (W.PS), and covariate balance under different weighting methods for White–Asian and White–Hispanic comparisons. In the White–Asian comparison, MEPS’s oversampling strategy leads to highly concentrated survey weights for Asians and a broader range with extreme weights for Whites, amplifying disparities in baseline covariates like BMI. These imbalances are reflected in the propensity score distributions, estimated using the survey-weighted propensity score model (W.PS), which show limited overlap between the groups, with most Whites concentrated at high scores (above 0.9) and Asians clustering near 0.8–0.9. The love plots further demonstrate that this limited overlap results in poor covariate balance under PATE and PATT, particularly under C.PS, where many covariates exceed the 0.1 PSMD threshold. Conversely, overlap weighting (PATO), especially under W.PS, achieves exact balance by focusing on comparable subpopulations, mitigating instability caused by extreme weights. In the White–Hispanic comparison, stronger overlap and more balanced survey weights yield better covariate alignment across methods, with PATO consistently outperforming other estimands in achieving robust and interpretable results. Overall, these findings underscore the importance of selecting appropriate weighting methods. Poor balance under PATE and PATT, especially in White–Asian comparisons, will inflate future variance estimation and bias effect estimates due to skewed target populations and extreme weights. Overlap weighting for PATO, particularly under W.PS, emerges as the most effective strategy for achieving population-level covariate balance and robust inference by focusing on comparable subpopulations and minimizing instability from extreme scores.[36, 45, 40, 42]
Table 8: Estimated treatment effects and standard errors (SE) for PATE, PATT, and PATO using different estimators under W.PS in the White–Asian and White–Hispanic comparisons of annual healthcare expenditures (2009 USD), in the MEPS study.
Comparison Group PSW MOM CVR WET PATE (SE) PATT (SE) PATO (SE)
White–Asian 2263.228 (575.601) 2327.342 (600.877) 1352.879 (236.564)
3962.495 (1913.583) 4053.586 (1958.159) 1349.055 (236.329)
1237.268 (614.993) 1231.726 (652.689) 1497.991 (283.749)
1179.610 (424.043) 1152.960 (442.362) 1358.156 (235.356)
White–Hispanic 607.782 (390.542) 463.373 (447.725) 1317.004 (164.707)
1230.326 (311.024) 1206.963 (352.923) 1308.903 (161.845)
1157.681 (317.028) 1172.387 (365.968) 1288.049 (164.289)
1242.545 (241.151) 1227.124 (264.413) 1308.268 (161.411)
Table 8 summarizes the estimated differences in annual healthcare expenditures (in 2009 USD) between Whites and Asians and between Whites and Hispanics for PATE, PATT, and PATO under W.PS. In the White–Asian comparison, limited overlap, particularly among high-BMI Asians, resulted in a PATE estimate using PSW of \$2263.23 (SE = 575.60), while the PATO estimate, focusing on comparable overlap subpopulations, had a notably smaller SE of 236.56 with a point estimate of \$1352.88, highlighting the stability of overlap-based estimands. Augmented estimators, especially WET, maintained consistent performance and resilience to extreme weights, producing a PATO estimate of \$1358.16 (SE = 237.55). Similarly, for the White–Hispanic comparison, the PATE estimate under PSW was \$607.78 (SE = 390.54), and the PATO estimate achieved a smaller SE of 164.71 with a point estimate of \$1317.00. Augmented estimators further reduced variability, with WET yielding a PATO estimate of \$1308.27 (SE = 162.38), reflecting robust performance across different racial comparisons under both favorable and challenging overlap conditions. Overall, these results underscore the effectiveness of W.PS in addressing survey weight imbalances across minority groups and demonstrate the robustness of WET in achieving precise treatment effect estimation, emphasizing their utility for analyzing racial disparities in MEPS. [36, 45]

### v2

@@ 4.2.2 Is there a global winner among the augmented estimators when all models are correctly specified?
Our simulation results, when both the propensity score and outcome models are correctly specified, indicate that all three augmented estimators consistently outperform PSW alone. While CVR demonstrate high efficiency under good overlap (Table 5), it can be sensitive to different sampling mechanisms and overlap scenarios. MOM offers consistent robustness but provides less bias control and coverage stability compared to WET across all estimands. Consequently, WET with W.PS emerges as the most reliable and efficient estimator in finite samples, maintaining near-nominal coverage, lower bias, particularly under poor overlap (Table 6) and treatment-dependent sampling (Web Table 9 and 10).
@@
To be more explicit, we first focus on the ”Cor|Cor” columns under good overlap (Table 5), W.PS-based MOM, CVR, and WET outperform PSW across all estimands (PATE, PATT, and PATO). The CVR estimator achieves significant efficiency gains, notably for PATT and PATO. Meanwhile, WET consistently performs slightly better than MOM, maintaining around 93% coverage and biases below 2% across most rows.
@@
Under poor overlap (”Cor|Cor” columns in Table 6) PSW alone produces noticeable biases (e.g., 8.848% for PATE) and reduced coverage (76.1%). This highlights the well-known vulnerability of simple weighting methods under poor overlap.[31] In contrast, MOM, CVR, and WET demonstrate lower biases and improved efficiency and coverage. For example, CVR with W.PS achieves nearly twenty-fold efficiency gains for PATE, reaching a relative efficiency of 23.8. Similarly, CVR boosts efficiency for PATT and PATO while maintaining better control over coverage. However, comparing across the augmented estimators, CVR and MOM show greater sensitivity under poor overlap. For instance, CVR achieves only 80% coverage for PATE, and MOM drops to 57.8% coverage for PATT. In comparison, WET consistently maintains coverage near or above 90%. Among these, WET with W.PS stands out, achieving the highest nominal coverage for PATO at 94.2% and outperforming other estimators.
@@
When sampling probabilities depend on treatment assignment (Web Tables 9 and 10), the advantage of WET becomes more evident. Under good overlap scenarios (Web Table 9), WET with W.PS achieves minimal biases and relative efficiency exceeding 1.5 for PATE and PATT while maintaining nominal coverage around 94%. By contrast, CVR suffers from larger bias and under-coverage, particularly for PATT. Similar patterns are found under poor overlap (Web Table 10).
@@
Taken together, WET with W.PS excels under complex sampling mechanisms. The weighted regression framework effectively minimizes bias and maintains stable coverage, even under challenging scenarios with poor overlap. While MOM provides consistent robustness and CVR achieves high efficiency under ideal conditions as in Table 5, both tend to be unstable when survey weights are increasingly heterogeneous. Thus, WET with W.PS stands out as a robust choice in our empirical investigation, particularly when the sampling mechanism depends on treatment and generates highly variable survey weights.
@@
4.2.3 Under misspecification, which methods are more robust and should be recommended under the complex survey setting?
@@
When model specifications deviate from the truth, the performance of estimators varies by estimand and overlap. When the propensity score model is misspecified but the outcome model is correct under good overlap (“Mis|Cor” columns in Table 5), all three augmented estimators achieve nominal coverage for PATE, demonstrating their doubly robust properties compared to PSW alone, which struggles to achieve even 80% coverage. MOM and WET control bias and coverage better than CVR. In poor overlap scenarios (Table 6), CVR achieves smaller bias but suffers reduced coverage (80%). In contrast, WET with W.PS maintains robustness, achieving 94.1% coverage for PATO, even with misspecified propensity score models.
@@
When the propensity score model is correct but the outcome model is misspecified (“Cor|Mis” columns), CVR demonstrates notable instability. Its bias increases dramatically for PATE under good overlap (Table 5), reaching nearly 10% compared to 0.2% bias for MOM and WET, and its coverage drops below 50% under poor overlap (Table 6). This empirical finding highlights the reliance of CVR on the correct outcome model to manage the extreme values introduced by clever covariates in poor overlap scenarios. In contrast, WET consistently performs best, achieving the optimal efficiency-coverage balance for PATO, despite the poor overlap and outcome model misspecification.
@@
Under the most challenging scenario, where both the propensity score and outcome models are misspecified (“Mis|Mis” columns), all methods experience significant biases and deteriorated coverage. Despite these challenges, WET remains the most reliable option, consistently maintaining coverage around 80–85% under good overlap, while CVR and MOM frequently see coverage drop below 80%. The extent of bias varies depending on the type of misspecification. In our simulation settings, where interaction terms are omitted from both models, this misspecification strongly affects propensity score estimation and outcome regression, leading to more biased estimates and poor coverage across all estimators.
@@
In conclusion, while CVR demonstrates potential for high efficiency with correctly specified models under good overlap, its heightened sensitivity to model misspecification and poor overlap may limit its practical reliability. In contrast, both MOM and WET exhibit greater robustness, with WET emerging as the most resilient choice across varying simulation settings.
5 Two case studies
@@ 5.1 Application to the ECLS-K dataset
In this first data example, we present a real-world example with multi-stage sampling to demonstrate the application of the different methods for using propensity score weighting with complex survey data. The data used are from the Early Childhood Longitudinal Study, Kindergarten class 1998-1999 (ECLS-K), [32] a nationally representative longitudinal study that tracks early school experiences of a large cohort from kindergarten through eighth grade in the United States. The study collects comprehensive information at the child, household, and school levels, with survey weights reflecting the multi-stage probability sampling design employed to ensure representativeness. [33] The participants were selected prior to the assignment of treatment. Specifically, children were sampled at the start of the study upon their entry into kindergarten, and subsequent treatment, such as the receipt of special education services, was determined in later stages, typically based on administrative records collected during follow-up data collection waves, such as in first grade or fifth grade. This example reflects a prospective study design, consistent with our main simulation setting, in which the sampling process does not depend on treatment assignment and survey weights are explicitly known. This design simplifies the integration of survey weights into subsequent estimations by reducing the steps required. Corresponding estimators tailored for prospective studies are further detailed in the appendix. The treatment status was defined as participation in special education services, as documented in school administrative records. The outcome of interest was the IRT-scaled math achievement test score from the ECLS-K dataset. A set of 39 baseline covariates covering demographic, household, and school characteristics were included to address potential confounding, as described in Keller and Tipton.[34]
@@
Web Figure 1 illustrates the population-level covariate balance, measured by Population-level survey-weighted Standardized Mean Differences (PSMD), which align with the weights $(\omega_{1,p}(\mathbf{X}_{i}),\omega_{0,p}(\mathbf{X}_{i}))$ used in the effect estimation and correspond to the specifications outlined in Table 1 for each target population and estimand. The PSMD for a baseline covariate $X$ is calculated as the absolute difference between the weighted means of $X$ for the treated and control groups, scaled by the pooled weighted standard deviation $s_{p}$. Specifically, PSMD is defined as $|\bar{X}_{1}-\bar{X}_{0}|/s_{p}$, where the weighted means are $\bar{X}_{1}=\left(\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})X_{i}Z_{i}\right)/\left(\sum_{i=1}^{n}\omega_{1,p}(\mathbf{X}_{i})Z_{i}\right)$ for the treated group and $\bar{X}_{0}=\left(\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})X_{i}(1-Z_{i})\right)/\left(\sum_{i=1}^{n}\omega_{0,p}(\mathbf{X}_{i})(1-Z_{i})\right)$ for the control group. The pooled weighted standard deviation, $s_{p}$, reflects a composite measure of the variability of $X$ within the treated and control groups, weighted according to the propensity score weights $\omega_{1,p}$ and $\omega_{0,p}$, and accounts for the sample size in each group. The figure compares the covariate balance for each population-level estimand—PATE, PATT, and PATO—using survey-weighted regression (W.PS). The visual evidence clearly demonstrates that W.PS achieves superior covariate balance, as indicated by a larger proportion of covariates achieving PSMDs below the 0.10 threshold, a conventional threshold for adequate balance.[35] This is further detailed in the grey-shaded cells of Web Table 17, which shows how W.PS systematically improves covariate balance across most baseline characteristics. Specifically, W.PS achieves exact balance for PATO, as a result of Proposition 1. We then estimate the average causal effect of elementary school special education services on fifth-grade math achievement, and the results are presented in Figure 1(a), with numerical details provided in Web Table 27.
(a) ECLS-K: Effect of special education on math achievement
(b) MEPS: White–Asian comparison
(c) MEPS: White–Hispanic comparison
Figure 1: Estimated treatment effects with 95% confidence intervals across estimators (PSW, MOM, CVR, WET) and estimands (PATE, PATT, PATO). Panel (a) shows the effect of special education services on fifth-grade math achievement in the ECLS-K study. Panels (b) and (c) show racial disparities in annual healthcare expenditures (2009 USD) in the MEPS study. The horizontal line indicates zero effect or disparity.
Across the three estimands targeting distinct populations, the results reveal notable differences in precision and stability. The PATE estimates the average effect across the entire student population, yielding negative effects ranging from 3.31-3.31 to 4.86-4.86 depending on the estimator, with PSW showing particularly large standard errors (SE =2.64=2.64). The PATT focuses on students who actually received special education services, where most of the augmented estimators (MOM, CVR, WET) yield statistically significant estimates, while PSW (1.12-1.12, SE =2.80=2.80) fails to detect the effect and produces a much wider confidence interval including zero. The PATO targets students for whom the special education placement decision is most uncertain and comparative evidence is most needed. For this estimand, all estimators produce more similar point estimates (ranging from 4.65-4.65 to 5.96-5.96) with comparable or smaller standard errors. For PATO, the WET estimate indicates that among these comparable students, those receiving special education scored approximately 5.8 points lower in math. Interestingly, among the augmented estimators, WET demonstrates particularly stable performance across all three estimands. These patterns align with our simulation observations.
@@ 5.2 Application to the 2009 Medical Expenditure Panel Survey dataset
In our second example, we demonstrate how the proposed methods perform in a retrospective survey design where the sampling mechanism depends directly on the exposure, specifically race. We use data from the 2009 Medical Expenditure Panel Survey (MEPS),[36] which historically oversamples minority groups, including Hispanics and Blacks, to improve the precision of health disparity analyses. Starting in 2006, this strategy expanded to include Asian households in response to changing demographics and analytical priorities.[37] This oversampling design ensures adequate representation of minority groups but can also lead to disproportionate survey weights, particularly for subgroups with limited population overlap. Unlike Data Example 1 (ECLS-K), where participants were sampled before treatment status was realized, MEPS employs a retrospective design where individuals are sampled after their race is determined, leading to survey weights that can correlate with race and demographic strata.
@@
The 2009 MEPS public-release file includes 9830 non-Hispanic White adults (hereafter “Whites”), 4020 Blacks, 1446 Asians, and 5150 Hispanics, each assigned a survey weight reflecting their probability of selection and nonresponse adjustments.[38] Here, we focus on two pairwise comparisons (White–Asian and White–Hispanic) to estimate racial disparities in total annual healthcare expenditures, measured in 2009 USD. Although “race” is not a manipulable treatment, the methodological goal parallels standard propensity score approaches to balancing covariates between two groups. Following the framework for unconfounded descriptive comparisons described in Li and Li,[39] we control for a set of variables reflecting clinical need, as instructed by the Institute of Medicine (IOM),[40] including body mass index (BMI), SF-12 physical and mental component summaries, health conditions, age, gender, and marital status. Additionally, we account for selected socioeconomic and functional limitation variables, consistent with prior applications to the MEPS data without adjusting for survey weights.[39]
@@
Web Figures 2, 3, and 4 collectively illustrate the interplay between survey weights, propensity score distributions estimated by the survey-weighted propensity score model (W.PS), and covariate balance under different weighting methods for White–Asian and White–Hispanic comparisons. In the White–Asian comparison, MEPS’s oversampling strategy leads to highly concentrated survey weights for Asians and a broader range with extreme weights for Whites, amplifying disparities in baseline covariates like BMI. These imbalances are reflected in the propensity score distributions, estimated using the survey-weighted propensity score model (W.PS), which show limited overlap between the groups, with most Whites concentrated at high scores (above 0.9) and Asians clustering near 0.8–0.9. The love plots further demonstrate that this limited overlap results in poor covariate balance under PATE and PATT, particularly under C.PS, where many covariates exceed the 0.1 PSMD threshold. Conversely, overlap weighting (PATO), especially under W.PS, achieves exact balance by focusing on comparable subpopulations, mitigating instability caused by extreme weights. In the White–Hispanic comparison, stronger overlap and more balanced survey weights yield better covariate alignment across methods, with PATO consistently outperforming other estimands in balancing covariates in the target population.
Figure 1(b) and (c) display the estimated differences in annual healthcare expenditures (in 2009 USD) between Whites and Asians and between Whites and Hispanics, with numerical details provided in Web Table 28. In the White–Asian comparison, limited overlap leads to substantial variability under PSW, with PATE and PATT estimates showing large standard errors (SE =575.60=575.60 and $600.88$, respectively). The augmented estimators improve precision considerably, with WET yielding a PATE estimate of \$1179.61 (SE =424.04=424.04). The PATO emphasizes a naturally comparable subpopulation with similar health status that is well-represented among both Whites and Asians. For this estimand, all estimators produce more similar estimates (ranging from \$1319.22 to \$1497.99) with consistently smaller standard errors around 236 to 284. In the White–Hispanic comparison, stronger overlap yields better precision overall compared to the White–Asian comparison. However, PSW estimates for PATE (\$607.78, SE =390.54=390.54) and PATT (\$463.37, SE =447.72=447.72) still show notable variability, while PATO estimates cluster tightly around \$1288 to \$1317 with standard errors near 161 to 165. Among the augmented estimators, WET again achieves the smallest standard errors across both comparisons, yielding PATO estimates of \$1358.16 (SE =236.36=236.36) for White–Asian and \$1308.27 (SE =161.41=161.41) for White–Hispanic, matching the observations in our simulation studies.